OpenStudy (anonymous):

Monica's heart beat 225 times in 3 minutes. Jordy's heart beat 360 times in 5 minutes. Whose heart beat at the faster rate and what was that rate? A. Monica had the faster heart rate of 72 beats per minute. B. Monica had the faster heart rate of 75 beats per minute. C. Jordy had the faster heart rate of 72 beats per minute. D. Jordy had the faster heart rate of 75 beats per minute.

OpenStudy (oaktree):

To do this, we can divide both of the numbers of beats by the number of minutes to get the rate per ONE minute, which makes it easier to compare. Can you do that?

OpenStudy (campbell_st):

find there heart beat per minute Monica = 225/3 jordy = 360/5 now compare the answers...

OpenStudy (anonymous):

idk?

OpenStudy (oaktree):

Like campbell_st said, just divide 225 by 3 and 360 by 5. What do you get?

OpenStudy (anonymous):

75/72

OpenStudy (campbell_st):

so monica's heart rate is 75 beats per minute jordy's heart rate is 72 beats per minute... that's the vital information... so who has the faster.. heart rate..?